7. How many molecules are present in 0.340g of HCI? A. 1.78 X 10-22 B. 6.46 X 1025 C. 5.61 X 1021 D. 2.06 X 10-23
Answer #1


Molar mass of HCl,
MM = 1*MM(H) + 1*MM(Cl)
= 1*1.008 + 1*35.45
= 36.458 g/mol


mass(HCl)= 0.340 g

use:
number of mol of HCl,
n = mass of HCl/molar mass of HCl
=(0.34 g)/(36.46 g/mol)
= 9.326*10^-3 mol
use:
number of molecules = number of mol * Avogadro’s number
number of molecules = 9.326*10^-3 * 6.022*10^23 molecules
number of molecules = 5.61*10^21 molecules
Answer: C